# Towards SLO-Optimized LLM Serving via Automatic Inference Engine Tuning

### v1

@@ V-C Solution Quality
For multi-objective optimization, we assess the quality of solutions from two perspectives. First, the optimality of solutions, whether optimal trade-offs have been achieved between TTFT and TPOT, that is, whether the optimized parameter configurations lie on the Pareto frontier of TTFT and TPOT. Second, we focus on the diversity of solutions, that is, whether the range of optimized parameter configurations on the Pareto frontier is sufficiently broad to offer customers more optional promising configurations. A high diversity of solutions can enhance the flexibility of SLO optimization.
@@
In Fig. 14 and Fig. 15, we plot the Pareto sets of configuration points sampled in the process of tuning the BOT and SQL applications under 4A10 for SCOOT and baselines. The plotted points represent the “local optimal solutions” identified by each tuning method. Additionally, we mark the global Pareto set from these local optimal configuration points with $\boldsymbol{\star}$. From Fig. 14, we can observe that for BOT, the majority of global optimal configuration points are contributed by SCOOT, both under LLAMA-13B and LLAMA-7B. Additionally, SCOOT yields a wider range of these global optimal points compared with baselines. From Fig. 15, we can see that for SQL, all the global optimal configuration points are contributed by SCOOT under both LLAMA-13B and LLAMA-7B. These experimental results confirm the superiority of SCOOT in both the optimality and diversity of SLO optimization in multi-objective situations.
(a) Configs under LLAMA2-7B.
(b) Configs under LLAMA2-13B.
Figure 14: Solution quality for BOT.
(a) Configs under LLAMA2-7B.
(b) Configs under LLAMA2-13B.
Figure 15: Solution quality for SQL.
(a) SLO optimization for CLS.
(b) SLO optimization for BOT.
(c) Performance tuning time for BOT and SQL.
Figure 16: Tuning efficiency of various parallelism degrees (PDs).
V-D Tuning Efficiency
In this subsection, we evaluate the tuning efficiency for SCOOT from two aspects. First, we focus on the speedup brought by the parallel suggestion. Second, we evaluate the efficiency of BO algorithms in making suggestions. We conduct experiments to tune the CSL and BOT applications under 2A100 with different parallelism degrees. In the experiments, the number of observations is also limited to 30 when the parallelism degree is set to 2. Figures 16(a) and 16(b) show the SCOOT’s performance in SLO optimization under various parallelism degrees, while Fig. 16(c) presents the total tuning time across various parallelism degrees.
From Figures 16(a) and 16(b), we can observe that the parallel suggestion achieves the same performance in SLO optimization as the non-parallel suggestion when the parallelism degree is set to 2. Additionally, as seen in Fig. 16(c), when setting the parallelism degree to 2, the parallel suggestion technique significantly reduces the total tuning time, nearly cutting it in half, which demonstrates that parallel suggestions can effectively improve tuning efficiency. Furthermore, from Fig. 16(c), we can also find that compared with the time of configuration evaluation, the time required for configuration suggestion is negligible, which validates the efficiency of applying BO algorithms to tune LLM inference engines. Moreover, the time for configuration suggestion also drastically decreases with the parallel suggestion. This is because, for both multi-objective and single-objective optimization, the parallel suggestion technique only samples additional points in the final sampling step, incurring negligible computational overhead. Since the number of recommended points is doubled, with the total number of observations fixed, the number of configuration suggestions is reduced to half, resulting in the time required for suggesting configuration points being cut in half as well. Compared to baselines, SCOOT can utilize more computational resources to accelerate the tuning process using the parallel suggestion, offering greater flexibility in efficiency improvement.
We also increase the parallelism degree and conduct experiments. However, we find that when the parallelism degree is set to a large value, such as greater than 4, SCOOT’s performance of SLO optimization is compromised. This is because when the parallelism degree increases, the surrogate model in each configuration suggestion is not as accurate as when the parallelism degree (PD) is set small. Let’s consider the following scenario. During the $i$th tuning iteration, the size of the observation set is $n$. For a case where PD=$k(k>1)$, SCOOT suggests $k$ configurations in one go, denoted as the $n+1$th, $n+2$th,…, $n+k$th configuration. All the configurations are suggested based on the surrogate model built with $n$ observations. When PD is small, for example, PD=1, the $n+1$th configuration is suggested based on the previous $n$ observations, but the $n+2$th configuration is suggested based on $n+1$ observations, making the prediction more accurate.
Since the parallel suggestion can reduce the total tuning time by PD times when the number of observations is fixed, a relatively small PD setting that does not affect the optimization results can significantly improve tuning efficiency. In addition, we also experimentally find that when the PD is set to a large value, the problem of performance degradation in SLO optimization can be resolved by appropriately increasing the number of observations. In the future, we will explore the best number of observations under different parallelism degrees to find the most cost-effective acceleration solution.
VI Related Works
VI-A Advanced LLM Inference Techniques
Existing inference engines such as vLLM and TensorRT-LLM support a variety of advanced inference techniques to improve the speed and throughput of LLM inference, including a series of kernel-level optimizations such as continuous batching [6], paged attention [7], flash attention [32, 33], and chunked prefill [8]. Besides, prefix caching [34] is employed by inference engines to reuse computed key and value tensors of requests’ common prefix texts to reduce redundant memory allocation and computation for newly arrived requests, which can significantly improve the serving efficiency when requests share a long common prefix. Moreover, speculative decoding [17, 35, 36, 37, 38] is also adopted by inference engines to accelerate LLM inference, where an efficient draft model is leveraged to generate tokens, and the LLM occasionally refines the draft.
VI-B Automatic Performance Tuning Approaches
Performance tuning techniques have been widely applied across a wide variety of fields, including database tuning [39, 40, 41], Spark configuration tuning [42, 43, 44], and compiler optimization [25]. The vast majority of these performance tuning studies adopt BO to find optimized parameter configurations since BO is theoretically grounded and can efficiently learn the relationship between performance and parameters from evaluations, thus intelligently tuning parameters for performance improvement. Because of these advantages, BO is also employed in resource allocation studies [45, 46].
However, to the best of our knowledge, previous studies have not focused on tuning the parameters of LLM inference engines and can not address the three unique challenges faced in LLM inference engine tuning.
VII Conclusion and Future Work
In this paper, we propose SCOOT, a BO-based automatic inference engine tuning system, to optimize SLOs for LLM inference services. SCOOT leverages both SOBO and MOBO to handle various optimization objectives and utilizes random forest regression to learn hidden constraints during the tuning process. Additionally, SCOOT exploits the parallel suggestion to accelerate the tuning process using additional computing resources. Experimental results show that SCOOT can effectively speed up the tuning process and significantly optimize SLOs, improving the request throughput by up to 68.3%, reducing the request tail latency by up to 40.6%, and reducing the TTFT and TPOT by up to 99.8% and 61.0%, respectively, compared to the default parameter configuration and existing tuning methods. In the future, we intend to expand SCOOT’s capabilities to accommodate more inference engines, such as TensorRT-LLM.

### v2

@@ A.5. Solution Quality Experiments
For multi-objective optimization, we evaluate the quality of solutions from two perspectives. First, the optimality of solutions, that is, whether the optimized parameter configurations lie on the Pareto frontier of TTFT and TPOT. Second, we focus on the diversity of solutions, that is, whether the range of optimized parameter configurations on the Pareto frontier is sufficiently broad to offer customers promising optional configurations to select based on their TTFT and TPOT requirements.
@@
Figures 15 and 16 plot the Pareto sets of configuration points sampled in the process of tuning the BOT and SQL applications under 4A10 for SCOOT and baselines. The plotted points represent the “locally optimal solutions” identified by each tuning method. Additionally, we mark the global Pareto set from these locally optimal configuration points with $\boldsymbol{\star}$. From Fig. 15, we can observe that for BOT, the majority of globally optimal configuration points are contributed by SCOOT. Besides, SCOOT yields a wider range of its globally optimal points compared with baselines. From Fig. 16, we can find that for SQL, all the globally optimal configuration points are contributed by SCOOT. These experimental results confirm the superiority of SCOOT in both the optimality and diversity of SLO optimization in multi-objective situations.
SCOOT achieves a broader range of solutions because it uses EHVI as the acquisition function for multi-objective problems to optimize the Pareto front of TTFT and TPOT. Besides, since the default configuration is chosen as the reference point, EHVI ensures that the TTFT and TPOT of the configuration with maximal HV are both better than the default, achieving the goal of optimizing both objectives simultaneously. Since VBO also employs EHVI, its solution range is broad. However, because it cannot effectively handle constraints, the optimality of the solutions is compromised.
A.6. Real-World Performance Experiments
Two real-world LLM inference services at Ant Group aim to maximize throughput while maintaining acceptable TTFT and TPOT. Both services are applications for short story generation, with average input lengths of 1,000 and 500 tokens, respectively, and an average output length of 300 tokens for both. We refer to these two inference services as story-teller-long (STL) and story-teller-short (STS) separately. Both services deploy an LLM with 80 billion parameters on 8 NVIDIA L20 48GB GPUs. During the online serving of STS and STL, the client uses concurrency to limit the maximum number of requests the service handles simultaneously, and we also include the concurrency as a parameter for optimization.
We adopt vLLM-0.5.5 and TensorRT-LLM-0.15.0 as inference engines, respectively, and conduct experiments to optimize throughput for these two inference services. The TensorRT-LLM parameters to be tuned are listed in Table 3. We have compared the tuning results of SCOOT with the default parameters of the inference engine, the recommended parameters from human algorithm experts at Ant Group, and VBO. Since VBO has been previously demonstrated to always outperform both GA and RD strategies in our experiments, we have not included comparisons with RD and GA. We filter out parameter configurations that do not meet the requirements for TTFT and TPOT and select the configurations with the highest throughput. The experimental results are presented in Table 4, where the throughput is measured in requests per second (RPS).
Human experts always simply enable certain techniques, which leads to sub-optimal results. Both STL and STS requests have a short common prefix. For vLLM-0.5.5, since chunked prefill and prefix caching are incompatible, human experts enable chunked prefill for STL and prefix caching for STS, respectively. For TensorRT-LLM 0.15.0, human experts enable both chunked prefill and prefix caching for STL and STS. However, due to the lack of sophisticated tuning of the client concurrency and parameters such as tokens-per-block and max-batch-size, the throughput improvement is always limited. From the experimental results, we can see that SCOOT consistently improves the throughput of two real-world inference services under various inference engines and significantly outperforms all baselines, which confirms the universality and practicality of SCOOT.
Table 3. TensorRT-LLM PARAMETERS TO TUNE.
Configuration Parameter Type Range
tp-size Int [1, #GPUs]
max-batch-size Int [64, 8192]
max-num-tokens Int [64, 8192]
tokens-per-block Enum {4, 8, 16, 32, 64, 128}
use-paged-context-fmha Bool {True, False}
enable-chunked-context Bool {True, False}
enable-block-reuse Bool {True, False}
capacity-scheduler-policy Enum {FCFS, EQUAL}
context-chunking-policy Enum {MAX_UTI, NO_EVICT}
Table 4. THROUGHPUT OPTIMIZATION RESULTS FOR REAL-WORLD INFERENCE SERVICES.
vLLM TensorRT-LLM
Method STL STS STL STS
Default 0.527 0.948 0.557 0.985
Expert 0.528 1.004 0.598 1.039
VBO 0.595 1.010 0.670 1.180
SCOOT 0.695 1.211 0.763 1.300
A.7. Discussion
A.7.1. Application Scope
SCOOT only requires the LLM inference engine to provide tunable parameters and an API-enabled entry point (e.g., FastAPI or OpenAI server). Since both widely-used LLM inference engines in cloud data centers (e.g., vLLM and TensorRT-LLM) and those deployed on personal devices with constraint resources (e.g., llama.cpp [49]) meet these requirements, SCOOT can be universally applied across a wide range of LLM deployment scenarios.
A.7.2. Parameter Selection
The selection of parameters and their value ranges significantly impact the tuning results. We select the parameters that are generally applicable across most scenarios and are of primary concern to customers, as indicated in the official documentation of vLLM [50] and TensorRT-LLM [51]. Furthermore, to ensure a more comprehensive optimization process, we intentionally include broad value ranges of the configuration parameters for performance tuning.
A.7.3. Final Deployed Product
We leverage existing company infrastructure and develop supplemental modules to ensure smooth deployment of SCOOT. Specifically, we implement several additional modules around SCOOT, including a dataset generation module to extract the request traces from the logging system into our database and utilize scripts to convert them into SCOOT-compatible formats, and a configuration parser to parse user-defined tuning parameters, including parameter names and their ranges. If parameters-to-tune are not specified by users, a default set of parameters and ranges will also be made available to all users. Besides, a post-processing module is developed to store the suggested parameters in the database and present them to customers for selection after SCOOT finishes parameter tuning.